OpenStudy (thatonegirl_):

Calculus! Find the coordinates of the points on the curve where the tangents are vertical.

OpenStudy (thatonegirl_):

\[x^2-xy+y^2=9\]

OpenStudy (thatonegirl_):

\[\frac{ dy }{ dx }=\frac{ y-2x }{ 2y-x }\]

OpenStudy (3mar):

Do you mean by tangent is the first derivative?

OpenStudy (thatonegirl_):

No i found the derivative before. I gotta find the tangent line where its vertical of the first equation, if thats what you're asking.

OpenStudy (3mar):

So how can I help you,sister?

OpenStudy (thatonegirl_):

How do I find the y or x value when the slope is undefined..?

hartnn (hartnn):

what is the slope of a vertical line?

OpenStudy (3mar):

it is 1/0

hartnn (hartnn):

the first derivative (or slope of tangent) is undefined. for that, denominator will be = 0 2y -x = 0

OpenStudy (thatonegirl_):

yes

OpenStudy (thatonegirl_):

Yeah i got to that but then what? Solve for y then plug that in for y?

hartnn (hartnn):

yes x =2y plug that in x^2−xy+y^2=9

OpenStudy (thatonegirl_):

Ahh thank you so much <3